OpenStudy (anonymous):

WILL GIVE MEDAL!!! Given the following information, determine the explicit equation for the geometric sequence. f(1) = 8, common ratio, r = 2 Give your answer starting with f(x)=. Use x as the variable. Do not include spaces. If an exponent is needed, use the ^. For example, x^2 would be written x^2. Use parantheses if needed to include multiple parts as an exponent. For example, 3^{(x+1)} typed as 3^(x+1) shows an exponent of x+1. Use either a * or place the second item in () to show multiplication. Example answer: f(x)=3(2)^x

OpenStudy (anonymous):

I got f(x)=8(2)^x is this correct?

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

for \[f(x)=8\times 2^x\] you would have \(f(1)=16\)

OpenStudy (anonymous):

Oh, so what do I do?

OpenStudy (anonymous):

use 4 instead of 8

OpenStudy (anonymous):

Okay thank you!